Given the information below for Seger Corporation, compute the expected share price at the end of 2017 using price ratio analysis. Assume that the historical average growth rates will remain the same for 2017. (Do not round intermediate calculations. Round your answers to 2 decimal places.)Year 2011 2012 2013 2014 2015 2016 Price $ 55.10 $ 61.00 $ 59.70 $ 57.20 $ 78.70 $ 94.10 EPS 2.60 3.31 4.11 4.81 7.35 8.35 CFPS 7.62 8.52 8.92 10.47 12.01 13.24 SPS 26.20 31.20 30.60 34.10 45.30 53.30
Answer #1

Please see the table below. The linkage will help you understand the mathematics. For the sale of explanation for the year 2011, P/E ratio has been calculated as = Price in 2011 / EPS in 2011 = 55.10 / 2.60 = 21.19 and so on. Growth rate in EPS for 2012 = EPS of 2012 / EPS of 2011 - 1 = 3.31/2.60 - 1 = 27.31% and so on. The last column is the average of the values in previous columns.

Year Linkage 2011 2012 2013 2014 2015 2016 Average
Price P             55.10        61.00        59.70        57.20        78.70        94.10
EPS E                2.60           3.31           4.11           4.81           7.35           8.35
CFPS C                7.62           8.52           8.92        10.47        12.01        13.24
SPS S             26.20        31.20        30.60        34.10        45.30        53.30
P/E             21.19        18.43        14.53        11.89        10.71        11.27        14.67
P/C                7.23           7.16           6.69           5.46           6.55           7.11           6.70
P/S                2.10           1.96           1.95           1.68           1.74           1.77           1.86
Growth rates
EPS 27.31% 24.17% 17.03% 52.81% 13.61% 26.98%
CFPS 11.81% 4.69% 17.38% 14.71% 10.24% 11.77%
SPS 19.08% -1.92% 11.44% 32.84% 17.66% 15.82%

EPS in 2017 = E2017 = E2016 x (1 + average growth rate in EPS) = 8.35 x (1 + 26.98%) = 10.60

CFPS in 2017 = C2017 = C2016 x (1 + average growth rate in CFPS) = 13.24 x (1 + 11.77%) = 14.80

SPS in 2017 = S2017 = S2016 x (1 + average growth rate in SPS) = 53.30 x (1 + 15.82%) = 61.73

Hence, the expected share price at the end of 2017 using

  • P/E ratio analysis = Average P/E x E2017 = 14.67 x 10.60 = $ 155.54
  • P/C ratio analysis = Average P/C x C2017 = 6.70 x 14.80 = $ 99.16
  • P/S ratio = Average P/S x S2017 = 1.86 x 61.73 = $ 115.12